OpenStudy (anonymous):

anyone? What is the value for x in the proportion x over x plus ten equals seven over twelve?

OpenStudy (anonymous):

\[\frac{ x }{ x + 10 }=\frac{ 7 }{ 12 }\] Is this what I'm looking at?

OpenStudy (anonymous):

I can't really tell here let me -write it so you can understand

OpenStudy (anonymous):

it's actually x/x+10 = 7/12

OpenStudy (anonymous):

ya you got it correct it showed in in writting it i refreshed my page but yes can you solve please?

OpenStudy (anonymous):

x=14

OpenStudy (anonymous):

thank you so much ma' am

OpenStudy (anonymous):

No problem.